How many ml of 95% alcohol should be mixed with 30% alcohol to make 23 fluid ounces of a 40% solution?
Answer #1
Component unit value Amount Value
95% alcohol solution 0.95 x 0.95x
30% alcohol solution 0.30 y 0.30y
Mixture 0.40

680.191

=23 ounces

0.40*680.91

=272.364

Sum of values of two alcohol = Value of mixture

0.95x + 0.30y = 272.364 equation 1

Sum of amounts of each alcohol = Amount of mixture

x + y = 680.191 equation 2

Solving both the equation,

Multiplying equation 2 with 0.30y

0.30x + 0.30y = 204.057

Subtracting equation 2 from equation 1

0.95x + 0.30y = 272.364

-

0.30x + 0.30y = 204.057

On solving:

0.65x = 68.307

x = 105.0877

y = 680.191 - 105.0877 = 575.1033

105.0877 ml of 95% alcohol should be mixed with 30% alcohol to make 23 fluid ounces of a 40% solution.